OpenStudy (anonymous):

Complete the square. Rewrite y = x2 + 14x + 29 in general form.

OpenStudy (anonymous):

x^2 +14x = -29 x^2 + 14x + 49 = 20 (x+7)^2 = 20 y = (x+7)^2 -20